OpenStudy (anonymous):

First six shipments: 10, 16, 22, 28, 34, 40 A factory is producing a linear number of objects in each shipment. If x = the shipment number, which function can be used to determine how many objects are in that shipment?

OpenStudy (anonymous):

I am not good with function...but notice that each shipment is increased by 6

OpenStudy (anonymous):

I think it might be : f(x) = 6x + 4 check.. first shipment f(1) = 6(1) + 4 f(1) = 6 + 4 f(1) = 10 2nd shipment f(2) = 6(2) + 4 f(2) = 12 + 4 f(2) = 16 it seems to check out ok... I think the answer is : f(x) = 6x + 4